OpenStudy (anonymous):

Without graphing, determine the equation whose graph intersects the graph of –6x + 3y = 11 exactly once. i)–6x + 3y = 13 ii) –24x + 12y = 44 iii) –4x + 3y = 11

OpenStudy (anonymous):

note that, the equations are in the form Ax + By = C.

OpenStudy (anonymous):

try to get the ratio A/B in each of the equation, and write it in simplest form. if its equals to the ratio of A/B, on -6x+3y=11, thus, the line, representing that equation does not intersect -6x+3y=11.

OpenStudy (anonymous):

But if its different on the A/B on -6x+3y =11, does, its intersects with the line representing -6x+3y=11. :))